i need help solving the following word problem:
A wholesaler has 100kg of mixed nuts that sell for $4/kg. In order to make the price more attractive, she plans to mix in some cheaper nuts worth $3.20/kg. If the wholesaler wants to sell the mixture for $3.40/kg, how many kilograms of the cheaper nuts should be used?
I came up with an answer of 75kg. Is this correct?

Found 2 solutions by Theo, mananth:
Answer by Theo(13342)

has 100 kg of mixed nuts that sell for $4.00/kg.

let x = amount that sell for $3.20/kg.

100*4 + x*3.20 = (100+x)*3.40

simplify to get:
400 + 3.2*x = 3.4*x + 340
subtract 340 from both sides of the equation to get:
60 + 3.2*x = 3.4*x
subtract 3.2*x from both sides of the equation to get:
60 = .2*x
divide both sides of the equation by .2 to get:
x = 300

she would need to add 300 kg of nuts at $3.20/kg.

if she did this, the total cost would be 100 * $4.00 + 300 * $3.20 = $400.00 + $960 = $1360 and the total cost/kg would be equal to $1360 / 400 = $3.40/kg

the total kg of her new mixture is 100 + 300 = 400

the total price of her new mixture is 100*4 + 300*3.2 = 1360

the cost per kg of her new mixture is 1360/400 = 3.4

your answer should be 300 kg.





Answer by mananth(16949)

..........cost $----------------quantity
Mixed nuts 4 ----------------100
Nuts B 3.2 ----------------x
Total 3.4 ----------------100+x
.....
4*100+3.2*x=3.4 (100+x)
400+3.2 x=340+3.4x
3.2x-3.4x=340-400
-0.2x=-60
/-0.2
x=300 kg.$ 3.2 nuts